Grand Garden is a luxury hotel with 160 sultes. Its regular sulte rate is $230 per night per sulte. The hotels cost per night is $155 per sulte and consists of the following $ 54 101 Variable direct labor and materials cost Pixed cost $5,870,000/10 suites) 35 days] Tosal eos per night pezsuite The hotel manager recelved an offer to hold the local Bikers Club annual meeting at the hotel In March, which is the hotels low season with an occupancy rate of under 60%. The Bikers Club would reserve 40 suites for three nights if the hotel could offer a 60% discount, or a rate of $92 per night. The hotel manager is Inclined to reject the offer because the cost per sulte per night is $155. Required: Prepare an analysis of this offer for the hotel manager. Contribution margin analysis
Answer #1

Contribution Margin Analysis

Revenue from 40 suites(40*$92*3)

$11040

Less:

Variable cost for 40 suites (40*54*3)

$6480

Contribution

$4560

The hotel manager should not reject the offer because there is increment in contribution margin by $4560. If the total capacity of a particular unit is not utilized to 100% then excess capacity can be sold to third party at variable cost. It means that for excess capacity which is unsold, variable cost is only to be taken for decision making process. Fixed cost is irrelevant in this case.

In the given question, Excess capacity is Number of Suites * Unoccupancy rate = 160*(100%- 60%) = 160*40% = 64. It means that 64 suites are vacant and only 40 suites are demanded by Biker’s club from these 64 unoccupied suites. These 64 unoccupied suites can be given at variable cost to be at no profit or loss situation. Any amount received above variable cost for these 64 unoccupied suites will be considered as profit.